OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = –3 and a5 = –768?

OpenStudy (anonymous):

Answer is -6912. Here how it is: \[a7=r(a1)^{7-1}\] \[a5=r(a1)^{5-1}\] \[a7/a5=(a1)^{^{2}}\] \[a7=a5\times(-3)^{2} since a1 = -3 given\] \[a7=(-768)\times9 = -6912\]

OpenStudy (anonymous):

thats not one of the answers in the multiple choice @vamshids

OpenStudy (anonymous):

oh I am really sorry for this I will tell u the right answer just now

OpenStudy (anonymous):

its ok thanks

OpenStudy (anonymous):

is the answer -12,288?

OpenStudy (anonymous):

YES! thanks!

OpenStudy (anonymous):

do you know how it came?

OpenStudy (anonymous):

in any GP: \[t _{n}=a(r)^{n-1}\]

OpenStudy (anonymous):

tn is the nth term of the series.. so accordingly subsititute your a1 and a5 to get r and then find a7!! its simple

OpenStudy (anonymous):

that helps so much! THanks!

OpenStudy (anonymous):

welcome bye!